Solve log(x) + log(x-3)=1
Answer by stanbon(48535)

Solve log(x) + log(x-3)=1
log[x(x-3)]=1
x^2-3x=10^1
x^2-3x-10=0
(x-5)(x+2)=0
x=5 or x=-2
x-2 is an extraneous solution
because log(-2) does not exist.
Final Solution: x=5
Cheers,
Stan H.